x + 3y = 32
-3x + 2y = 3
When I try to solve for either x or y variable i get a fraction and/or mixed number and I cant solve it.
Thank you for your help
Chris

Answer by sudhanshu_kmr(1152)

x + 3y = 32
-3x + 2y = 3
multiply first equation by 3 and add to second equation
you will get, 11y = 99
=> y = 9
put this value of first equation,
x + 27 = 32
=> x = 32-27 = 5
so, x= 5, y = 9

Everyone is welcome to get free online help: sudhanshu.cochin@gmail.com